2. Y = e^-2x . Sin 3x show that

d^2y/dx^2 = -13y - 4 dy/dx


1
Expert's answer
2020-08-31T17:22:56-0400

2.

"y=e^{-2x}sin3x"

"y'=3e^{-2x}cos3x-2e^{-2x}sin3x"

"y''=-9e^{-2x}sin3x-6e^{-2x}cos3x-6e^{-2x}cos3x+4e^{-2x}sin3x="

"=-5e^{-2x}sin3x-12e^{-2x}cos3x"

Then:

"-13y - 4y'=-13e^{-2x}sin3x-4(3e^{-2x}cos3x-2e^{-2x}sin3x)="

"=-13e^{-2x}sin3x-12e^{-2x}cos3x+8e^{-2x}sin3x="

"=-5e^{-2x}sin3x-12e^{-2x}cos3x=y''"